OpenStudy (anonymous):

what is f(-3) if f(x) = the absolute value of 2x-1 -5

OpenStudy (jdoe0001):

\(\bf f(x) = |2x-1 -5|\qquad \qquad f(\color{red}{-3})= |2(\color{red}{-3})-1 -5|\)

OpenStudy (anonymous):

i got that part bbut i dont know what to do after that

OpenStudy (anonymous):

how do i get the solution

OpenStudy (jdoe0001):

well, absolute value expressions are ALWAYS positive, so :)

OpenStudy (anonymous):

so its either 2 or 0 because thats the only choices that are positive haha

OpenStudy (jdoe0001):

\(\bf f(x) = |2x-1 -5|\qquad \qquad f(-3)= |2(-3)-1 -5|\implies |-6-1-5|\\ f(-3) = |-12| \implies f(-3) = 12\)

OpenStudy (jdoe0001):

hmmm well. what's the absolute value expression? |2x| -1 -5? or |2x-1| -5?

OpenStudy (anonymous):

the second one but you replace x with -3

OpenStudy (jdoe0001):

\(\bf f(x) = |2x-1| -5\qquad \qquad f(-3)= |2(-3)-1| -5\\ \quad \\ f(-3)= |-6-1| -5\) what would that give you?

OpenStudy (anonymous):

see i dont know what to do after those steps:/

OpenStudy (anonymous):

|-3*2-1|-5 = 7-5=2

OpenStudy (jdoe0001):

\(\bf f(x) = |2x-1| -5\qquad \qquad f(-3)= |2(-3)-1| -5\\ \quad \\ \textit{absolute value expressions are always positive}\quad thus\\ f(-3)= |-6-1| -5\implies f(-3)= |\color{blue}{-7}| -5\implies f(-3)= \color{blue}{7} -5\)

OpenStudy (anonymous):

haaha